A magnetic tape storing information in binary form
Answer #1

P(T\E) = P(ET)/ P(E)
Where P(T\E) = the probability that given you read a 1 that this is a correct reading.
P(ET) = probability that you will select a 1 times the probability that it will be detected as a 1
P(E) = P(ET) + P(ET")
Where P(ET) is as defined above, and P(ET") is the probability of selecting a 0 times the probability it will be detected as a 1
P(ET) = (0.5)(0.85) / [(0.5)(0.85) + (0.5)(0.1)]
P(ET) = 0.425/0.475

P(ET) = 0.8947